OpenStudy (anonymous):

Add. 18 5/12, + 8 10/12, + 23 7/12, + 28 7/17 A. 77 B. 77 1/2 C. 78 1/2 D. 79

OpenStudy (anonymous):

\[\large \sf 18\frac{5}{12}+8\frac{10}{12}+23\frac{7}{12}+28\frac{7}{17}\] \[\ \large (18+8+23+28)+(\frac{5}{12}+\frac{10}{12}+\frac{7}{12}+\frac{7}{17})\]

OpenStudy (anonymous):

Is the answer D?

OpenStudy (anonymous):

It's the closest, although I got 79.25

OpenStudy (anonymous):

I guess if they wanted you to round, then it is